OpenStudy (anonymous):

factoring quadratic polynomials x^2+2x-3 y^2+y-12 g^2-10g+21

OpenStudy (anonymous):

x^2+2x-3 = (x +3)(x-1)

OpenStudy (anonymous):

y^2+y-12 = (y + 4)( y -3)

OpenStudy (anonymous):

g^2-10g+21 = ( g -7)( g -3)